OpenStudy (katherinesmith):

Write log24 + log23 as a single logarithm.

OpenStudy (katherinesmith):

\[\log _{2}4 + \log _{2}3\]

OpenStudy (anonymous):

use \[\log_b(A)+\log_b(B)=\log_b(AB)\]

OpenStudy (katherinesmith):

@satellite73 what numbers plug in for each letter?

jimthompson5910 (jim_thompson5910):

compare \[\large \log_{2}(4) + \log _{2}(3)\] to \[\large \log_b(A)+\log_b(B)\]

OpenStudy (katherinesmith):

\[\log _{2}(4) + \log _{2}(3) = \log _{2}(12)\]

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (katherinesmith):

now what

jimthompson5910 (jim_thompson5910):

you're done, its a single log

OpenStudy (katherinesmith):

wow

jimthompson5910 (jim_thompson5910):

you're overthinking this one

OpenStudy (katherinesmith):

math is insane

jimthompson5910 (jim_thompson5910):

you'll get the hang of it soon